OpenStudy (anonymous):

Subtract the following radicals: 5√45 - 3√5 A. 5√5 B. 12 C. 12√5 D. 42√5

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (jdoe0001):

\(\large { 5\sqrt{45}-3\sqrt{5}\qquad {\color{brown}{ 45\to 3\cdot 3\cdot 5\to 3^2\cdot 5 }} \\ \quad \\ 5\sqrt{45}-3\sqrt{5}\implies 5\sqrt{{\color{brown}{ 3^2\cdot 5}}}-3\sqrt{5}\implies ? }\)

OpenStudy (anonymous):

How would I find that answer?

OpenStudy (jdoe0001):

take out what you can from the 1st radical what do you have left?

OpenStudy (anonymous):

9*5 is 45, so its 5 sqrt 45?

OpenStudy (jdoe0001):

well... the 3 comes out WITHOUT the exponent thus \(\large { 5\sqrt{45}-3\sqrt{5}\qquad {\color{brown}{ 45\to 3\cdot 3\cdot 5\to 3^2\cdot 5 }} \\ \quad \\ 5\sqrt{45}-3\sqrt{5}\implies 5\sqrt{3^2\cdot 5}-3\sqrt{5}\implies 5\cdot 3\sqrt{5}-3\sqrt{5} \\ \quad \\ 15{\color{blue}{ \sqrt{5} }}-3{\color{blue}{ \sqrt{5} }}\implies ? }\)

OpenStudy (anonymous):

Oh okay uhm its prime isnt it?

OpenStudy (jdoe0001):

the 5 is prime.... the 3 was squared notice the root is 2, so the 3 is raised at the "2" the root is at the "2" so the 3 comes out without its exponent but yes, the 5 is prime, stays inside

OpenStudy (anonymous):

Ok what next

OpenStudy (jdoe0001):

so if we made say \(\bf \sqrt{5} = a\) 15a - 3a = ?

OpenStudy (anonymous):

12

OpenStudy (jdoe0001):

yeap thus \(\large 15{\color{blue}{ \sqrt{5} }}-3{\color{blue}{ \sqrt{5} }}\implies 12{\color{blue}{ \sqrt{5}}}\)

OpenStudy (anonymous):

Thanks!